( )+( )+( )=30
Fill the ( ) use this (1,3,5,7,9,11,13,15)

Answer by farohw(175)


(9-1)+(11)+(11) = 30

(5)+(11-1)+(15) = 30

(1+5) + (9) + (15) = 30

When you add three odd numbers you will get an odd answer. So you have to apply other strategies to get an even answer of 30 using the odd numbers.